# DmC: Nearest Neighbor Guidance Diffusion Model for Offline Cross-domain Reinforcement Learning

### v1

@@ 1 Introduction
Previous research has tackled cross-domain RL in various settings, including online, where both domains are online [6, 27], or offline target with online source (hybrid) [34, 35]. In this work, we focus on the cross-domain offline setting [23, 24, 53], where both source and target domains are offline. This setting is crucial for enhancing the sample efficiency of offline RL methods [17, 16]. A key idea is to leverage source samples close to the target domain to enhance policy learning. Thus, a core challenge lies in accurately quantifying the domain (dynamics) gap between two domains. Existing works on cross-domain offline reinforcement learning (RL) have introduced various methods to measure the dynamics gap between domains, such as training domain classifiers [23], using conditional variational autoencoder (CVAE) models to approximate target dynamics [24], and estimating mutual information [48]. However, these approaches typically rely on large target datasets, which are impractical in many real-world applications, such as healthcare. To address this limitation, we focus on cross-domain offline RL in the limited target data setting, which brings several challenges that require novel solutions.
@@ 2.1 Cross-Domain Offline RL
Cross-domain RL aims to improve sample efficiency in the target domain by leveraging data from additional source environments. Several approaches have been proposed to address this challenge, including system identification [49, 54, 2], domain randomization [40, 44, 36], and meta-RL [7, 32, 51]. However, these methods often require environment models or domain knowledge to carefully select randomized parameters. Recently, several methods have attempted to measure dynamics discrepancy for various settings, including purely online [6, 18, 52], purely offline [24, 48, 53, 3], or a hybrid setting [34]. In this work, we focus on the cross-domain offline setting, where both the source and target domains are offline. Previous approaches tackle this problem through reward modification [23], support constraints [24], or data filtering via mutual information [48]. However, these methods require training neural networks (NNs) to estimate the dynamics gap between two domains, which can be challenging when only limited target data is available. To overcome this limitation, we propose using $k-$NN based estimation to measure the divergence between source and target domains, avoiding the need for NNs training. The concurrent work from Lyu et al. [30] introduced OTDF, which leverages optimal transport (OT) to address neural network training challenges. However, OTDF involves solving a computationally expensive OT problem between source and target datasets. In contrast, $k$-NN estimation offers a more efficient and scalable solution. Moreover, unlike prior works, including OTDF, our approach generates more samples close to the target domain, thus bringing more sample efficiency. We employ a novel guided-diffusion model to augment the source dataset with samples close to the target domain. These strategies enable our method to perform effectively in limited target data settings.
@@ 3.1 Reinforcement Learning
We address the cross-domain offline setting, where the source domain is defined as $\mathcal{M}_{src}=\left(\mathcal{S},\mathcal{A},\gamma,r,d_{0},P_{src}\right)$ and the target domain as $\mathcal{M}_{tar}=\left(\mathcal{S},\mathcal{A},\gamma,r,d_{0},P_{tar}\right)$. We assume that the two domains share the same state space $\mathcal{S}$, action space $\mathcal{A}$, discount factor $\gamma$, reward function $r$, and initial state distribution $d_{0}$, differing only in their dynamics models. In this setting, we have access to offline datasets collected from both domains: $D_{src}$ from the source domain and $D_{tar}$ from the target domain. The objective is to utilize the additional source dataset $D_{src}$ to improve policy learning for the target domain, using the target dataset $D_{tar}$. We further denote the empirical MDP induced by a dataset $D$ as ^$\widehat{\mathcal{M}}=\left(\mathcal{S},\mathcal{A},r,d_{0},\gamma,\widehat{P}\right)$, where $\widehat{P}$ represents the empirical transition dynamics derived from $D$. In the cross-domain offline setting, this results in two empirical MDPs: ^$\widehat{\mathcal{M}}_{src}$ with transitions $\widehat{P}_{src}$ and ^$\widehat{\mathcal{M}}_{tar}$ with transitions $\widehat{P}_{tar}$.
@@ 3.2 Diffusion Model
$q\left(x^{1:T}\mid x^{0}\right)&=\prod_{t=1}^{T}q\left(x^{t}\mid x^{t-1}\right)=\prod_{t=1}^{T}\mathcal{N}\left(\sqrt{\alpha^{t}}x^{t-1},\beta^{t}\mathrm{I}\right)\\ q\left(x^{t}\mid x^{0}\right)&=\mathcal{N}\left(\sqrt{\overline{\alpha}^{t}}x^{0},\left(1-\overline{\alpha}^{t}\right)\mathrm{I}\right),\;\text{with: }\overline{\alpha}^{t}=\prod_{i=1}^{t}\alpha^{i}\end{aligned}$
@@
[…]
@@ 5.1 $k$-NN Based Domain Gap Estimation
[…]
@@
$\widehat{\mathcal{D}}_{KL}(P_{src}||P_{tar})&=\frac{1}{N}\sum^{N}_{i=1}\left(\log\frac{M\nu^{d}_{k,\text{tar}}(i)}{(N-1)\nu^{d}_{k,\text{src}}(i)}\right)\\ &\propto 1/N\sum^{N}_{i=1}\left(\log(\nu^{d}_{k,\text{tar}}(i))-\log(\nu^{d}_{k,\text{src}}(i))\right)\end{aligned}$
@@
[…]
where $x_{src,i,k}$ and $x_{tar,i,k}$ are the $k$ nearest neighbors of $x_{src,i}$ in the source dataset $D_{src}$ and the target dataset $D_{tar}$ respectively. Intuitively, $\rho_{k}(x_{src,i})$ quantifies the discrepancy of the source sample $x_{src,i}$ to the target dataset, and is small if the distance from $x_{src,i}$ to its $k$-nearest neighbors in the target dataset is smaller than the distance to its $k$-nearest neighbors in the source dataset. The critical advantage of $k$-NN estimation lies in its consistency and simplicity: it is stable and avoids the need for neural network training, thus mitigating the risk of overfitting when estimating the domain gap. Furthermore, it is computationally efficient and easy to implement, leveraging the KD-tree data structure. In practice, we use FAISS library [5] in our implementation and compute all scores for 1 million source samples with 5000 target samples within 1 minute.
@@ 5.2 Nearest-Neighbor Guided Diffusion Model
[…]
@@
The training process for the diffusion model is done before the training of the RL policy. Once the diffusion model is trained, we generate source samples that are close to the target domain by setting the value of the conditional context $y$ to be higher than the top $\kappa\%$ of the source sample scores in source dataset $D_{src}$. Specifically, we uniformly sample a value $\chi$ from the range $[\kappa,100]$ and set $y$ to the $\chi$-th quantile of the source dataset’s score distribution. Finally, generated samples are combined with source dataset for later policy learning.
@@ 5.3 Algorithm
Based on the $k$-NN estimation as the domain gaps measurement and the guided-diffusion model for upsampling source dataset with generated samples close to the target domain, we obtain a practical policy adaptation algorithm, named DmC. We summarize our proposed algorithm in Algorithm 1. In practice, we normalize the value of $\rho_{k}$ to [0,1][0,1][ 0 , 1 ]. Specifically, we first subtract each value with the minimum value of $\rho_{k}$ in the source dataset $D_{src}$ to adjust the value to the range [0,+][0,$\infty$][ 0 , + ∞ ], and then scale it as follows:
@@
[…]
@@ 5.4 Theoretical Analysis
$&J^{P_{tar}}(\pi)-J^{\widehat{P}_{src}}(\pi)\\ &\geq-\frac{2r_{\text{max}}\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s,a\sim d^{\pi}_{P_{tar}},s^{\prime}\sim P_{tar}}\left[D_{TV}\left(\pi(.|s^{\prime})||\pi_{D_{tar}}(.|s^{\prime})\right)\right]\\ &-\frac{2r_{\text{max}}\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s,a\sim d^{\pi_{D_{tar}}}_{P_{tar}},s^{\prime}\sim P_{tar}}\left[D_{TV}\left(\pi(.|s^{\prime})||\pi_{D_{tar}}(.|s^{\prime})\right)\right]\\ &-\frac{2r_{\text{max}}\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s,a\sim d^{\pi_{D_{tar}}}_{P_{tar}}}\left[\sqrt{1/2D_{KL}(P_{tar}(.|s,a)||\widehat{P}_{src}(.|s,a))}\right].\end{aligned}$
@@
2:Initialize: Value function $Q$, policy $\pi$, diffusion model $G_{\theta}$, $k$ NN value, coefficient $\lambda$, source score buffer $SB$.
@@
7: Train $G_{\theta}$ with $D_{src}$ and $SB$ via Eq. (8).
8: Generate samples from the diffusion model and add to the source dataset $D_{src}$ and source score buffer $SB$.

### v2

@@ 1 Introduction
Previous research has tackled cross-domain RL in various settings, including online, where both domains are online [6, 27], or offline target with online source (hybrid) [34, 35, 58]. In this work, we focus on the cross-domain offline setting [23, 24, 53], where both source and target domains are offline. This setting is crucial for enhancing the sample efficiency of offline RL methods [17, 16, 59]. A key idea is to leverage source samples close to the target domain to enhance policy learning. Thus, a core challenge lies in accurately quantifying the domain (dynamics) gap between two domains. Existing works on cross-domain offline reinforcement learning (RL) have introduced various methods to measure the dynamics gap between domains, such as training domain classifiers [23], using conditional variational autoencoder (CVAE) models to approximate target dynamics [24], and estimating mutual information [48]. However, these approaches typically rely on large target datasets, which are impractical in many real-world applications, such as healthcare. To address this limitation, we focus on cross-domain offline RL in the limited target data setting, which brings several challenges that require novel solutions.
@@ 2.1 Cross-Domain Offline RL
Cross-domain RL [57, 61] aims to improve sample efficiency in the target domain by leveraging data from additional source environments. Several approaches have been proposed to address this challenge, including system identification [49, 54, 2], domain randomization [40, 44, 36], and meta-RL [7, 32, 51]. However, these methods often require environment models or domain knowledge to carefully select randomized parameters. Recently, several methods have attempted to measure dynamics discrepancy for various settings, including purely online [6, 18, 52, 60], purely offline [24, 48, 53, 3, 56, 55], or a hybrid setting [34]. In this work, we focus on the cross-domain offline setting, where both the source and target domains are offline. Previous approaches tackle this problem through reward modification [23], support constraints [24], or data filtering via mutual information [48]. However, these methods require training neural networks (NNs) to estimate the dynamics gap between two domains, which can be challenging when only limited target data is available. To overcome this limitation, we propose using $k$-NN based estimation to measure the divergence between source and target domains, avoiding the need for NNs training. The concurrent work from Lyu et al. [30] introduced OTDF, which leverages optimal transport (OT) to address neural network training challenges. However, OTDF involves solving a computationally expensive OT problem between source and target datasets. In contrast, $k$-NN estimation offers a more efficient and scalable solution. Moreover, unlike prior works, including OTDF, our approach generates more samples close to the target domain, thus bringing more sample efficiency. We employ a novel guided-diffusion model to augment the source dataset with samples close to the target domain. These strategies enable our method to perform effectively in limited target data settings.
@@ 3.1 Reinforcement Learning
[…]
@@ 3.2 Diffusion Model
$\begin{aligned} q\left(x^{1:T}\mid x^{0}\right)&=\prod_{t=1}^{T}q\left(x^{t}\mid x^{t-1}\right)=\prod_{t=1}^{T}\mathcal{N}\left(\sqrt{\alpha^{t}}x^{t-1},\beta^{t}\mathrm{I}\right)\\ q\left(x^{t}\mid x^{0}\right)&=\mathcal{N}\left(\sqrt{\overline{\alpha}^{t}}x^{0},\left(1-\overline{\alpha}^{t}\right)\mathrm{I}\right),\;\text{with: }\overline{\alpha}^{t}=\prod_{i=1}^{t}\alpha^{i}\end{aligned}$
@@
$\begin{aligned} q\left(x^{t-1}\mid x^{t},x^{0}\right)=\mathcal{N}\left(\frac{1}{\sqrt{\alpha^{t}}}\left(x^{t}-\frac{1-\alpha^{t}}{\sqrt{1-\overline{\alpha}^{t}}}\epsilon\right),\beta^{t}\mathrm{I}\right)\end{aligned}$
@@ 5.1 $k$-NN Based Domain Gap Estimation
$\begin{aligned} P_{D,k}(x^{i})=\frac{k/|D|}{\mathcal{V}(\mathcal{B}(x,\nu_{k,D}(x)))}=\frac{k}{|D|c\nu_{k,D}(x)},\end{aligned}$
@@
$\begin{aligned} \widehat{\mathcal{D}}_{KL}(P_{src}||P_{tar})&=\frac{1}{N}\sum^{N}_{i=1}\left(\log\frac{M\nu^{d}_{k,\text{tar}}(i)}{(N-1)\nu^{d}_{k,\text{src}}(i)}\right)\\ &\propto 1/N\sum^{N}_{i=1}\left(\log(\nu^{d}_{k,\text{tar}}(i))-\log(\nu^{d}_{k,\text{src}}(i))\right)\end{aligned}$
@@
$\begin{aligned} &\rho_{k}(x_{src,i})&=\log(\|x_{src,i}-x_{tar,i}^{k}\|_{2})-\log\left(\left\|x_{src,i}-x_{src,i}^{k}\right\|_{2}\right),\end{aligned}$
where $x_{src,i}^{k}$ and $x_{tar,i}^{k}$ are the $k$ nearest neighbors of $x_{src,i}$ in the source dataset $D_{src}$ and the target dataset $D_{tar}$ respectively. Intuitively, $\rho_{k}(x_{src,i})$ quantifies the discrepancy of the source sample $x_{src,i}$ to the target dataset, and is small if the distance from $x_{src,i}$ to its $k$-nearest neighbors in the target dataset is smaller than the distance to its $k$-nearest neighbors in the source dataset. The critical advantage of $k$-NN estimation lies in its consistency and simplicity: it is stable and avoids the need for neural network training, thus mitigating the risk of overfitting when estimating the domain gap. Furthermore, it is computationally efficient and easy to implement, leveraging the KD-tree data structure. In practice, we use FAISS library [5] in our implementation and compute all scores for 1 million source samples with 5000 target samples within 1 minute.
@@ 5.2 Nearest-Neighbor Guided Diffusion Model
$\begin{aligned} \min_{\theta}\mathbb{E}\left[\left\|x_{src,i}-G_{\theta}\left(x_{src,i}+\sigma^{t}\epsilon,m\cdot\rho_{k}(x_{src,i}),\sigma^{t}\right)\right\|^{2}\right].\end{aligned}$
@@
The training process for the diffusion model is done before the training of the RL policy. Once the diffusion model is trained, we generate source samples that are close to the target domain by setting the value of the conditional context $y$ to be higher than the top $\kappa\%$ of the source sample scores in source dataset $D_{src}$. Specifically, we uniformly sample a value $\chi$ from the range $\kappa,100]$ and set $y$ to the $\chi$-th quantile of the source dataset’s score distribution. Finally, generated samples are combined with source dataset for later policy learning.
@@ 5.3 Algorithm
Based on the $k$-NN estimation as the domain gaps measurement and the guided-diffusion model for upsampling source dataset with generated samples close to the target domain, we obtain a practical policy adaptation algorithm, named DmC. We summarize our proposed algorithm in Algorithm 1. In practice, we normalize the value of $\rho_{k}$ to $[0,1]$. Specifically, we first subtract each value with the minimum value of $\rho_{k}$ in the source dataset $D_{src}$ to adjust the value to the range $0,+\infty]$, and then scale it as follows:
@@
$\begin{aligned} \mathcal{L}_{Q}=\mathbb{E}_{D_{tar}}\left[(Q_{\phi}-\mathcal{T}Q_{\phi})^{2}\right]+\mathbb{E}_{D_{src}}\left[\omega(s,a,s^{\prime})(Q_{\phi}-\mathcal{T}Q_{\phi})^{2}\right].\end{aligned}$
@@ 5.4 Theoretical Analysis
$\begin{aligned} &J^{P_{tar}}(\pi)-J^{\widehat{P}_{src}}(\pi)\\ &\geq-\frac{2r_{\text{max}}\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s,a\sim d^{\pi}_{P_{tar}},s^{\prime}\sim P_{tar}}\left[D_{TV}\left(\pi(.|s^{\prime})||\pi_{D_{tar}}(.|s^{\prime})\right)\right]\\ &-\frac{2r_{\text{max}}\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s,a\sim d^{\pi_{D_{tar}}}_{P_{tar}},s^{\prime}\sim P_{tar}}\left[D_{TV}\left(\pi(.|s^{\prime})||\pi_{D_{tar}}(.|s^{\prime})\right)\right]\\ &-\frac{2r_{\text{max}}\gamma}{(1-\gamma)^{2}}\mathbb{E}_{s,a\sim d^{\pi_{D_{tar}}}_{P_{tar}}}\left[\sqrt{1/2D_{KL}(P_{tar}(.|s,a)||\widehat{P}_{src}(.|s,a))}\right].\end{aligned}$
@@
2:Initialize: Value function $Q$, policy $\pi$, diffusion model $G_{\theta}$, $k$ NN value, coefficient $\lambda$, source score buffer SBSB.
@@
7: Train $G_{\theta}$ with $D_{src}$ and SBSB via Eq. (8).
8: Generate samples from the diffusion model and add to the source dataset $D_{src}$ and source score buffer SBSB.